# A Characterization for Tightness of the Sparse Moment-SOS Hierarchy

### v1

@@ Acknowledgements
This project was begun at a SQuaRE at the American Institute of Mathematics (AIM). The authors would like to thank AIM for providing a supportive and mathematically rich environment. Jiawang Nie is partially supported by the NSF grant DMS-2110780. Zheng Qu is partially supported by the Hong Kong Research Grants Council GRF No. 17317122. Xindong Tang is partially supported by the Hong Kong Research Grants Council HKBU-15303423. Linghao Zhang is partially supported by the UCSD School of Physical Sciences Undergraduate Summer Research Award.
Appendix A Proof of Theorem 4.6
We first prove Theorem 4.6 for the case $m=2$.
Lemma A.1.
(A.1) $\left\{\begin{array}[]{cll}\displaystyle\min&f_{1}(x_{\Delta_{1}})\\ \mbox{s.t.}&h_{1}(x_{\Delta_{1}})=0,\\ &g_{1}(x_{\Delta_{1}})\geq 0,\end{array}\right.\quad\left\{\begin{array}[]{cll}\displaystyle\min&f_{2}(x_{\Delta_{2}})\\ \mbox{s.t.}&h_{2}(x_{\Delta_{2}})=0,\\ &g_{2}(x_{\Delta_{2}})\geq 0.\end{array}\right.$
(A.2) $f_{1}+p\geq 0\,\mbox{ on }\,K_{\Delta_{1}},\quad f_{2}-p-f_{\min}\geq 0\,\mbox{ on }\,K_{\Delta_{2}},$
Proof.
By the assumption, we can write that
(A.3) $U_{0}\coloneqq\mathfrak{p}_{12}(K_{\Delta_{1}})\cap\mathfrak{p}_{21}(K_{\Delta_{2}}),\enspace U_{1}\coloneqq\mathfrak{p}_{12}(K_{\Delta_{1}})\backslash U_{0},\enspace U_{2}\coloneqq\mathfrak{p}_{21}(K_{\Delta_{2}})\backslash U_{0}.$
Define the optimal value functions:
(A.6) $\left\{\begin{array}[]{rcl}F_{1}(u)\coloneqq&\displaystyle\min&f_{1}(x_{\Delta_{1}})\\ &\mbox{s.t.}&x_{\Delta_{1}}\in K_{\Delta_{1}},\,x_{\Delta_{12}}=u,\end{array}\right.$
(A.9) $\left\{\begin{array}[]{rcl}F_{2}(u)\coloneqq&\displaystyle\min&f_{2}(x_{\Delta_{2}})\\ &\mbox{s.t.}&x_{\Delta_{2}}\in K_{\Delta_{2}},\,x_{\Delta_{12}}=u.\end{array}\right.$
For each $u\in U_{0}\cup U_{1}$, the problem (A.6) is feasible and its optimal value is finite. Similarly, for each $u\in U_{0}\cup U_{2}$, the problem (A.9) is feasible and its optimal value is finite. For all $u\in U_{0}$, $F_{1}(u)+F_{2}(u)$ equals the minimum value of
$\left\{\begin{array}[]{cll}\displaystyle\min&f_{1}(x_{\Delta_{1}})+f_{2}(x_{\Delta_{2}})\\ \mbox{s.t.}&x_{\Delta_{1}}\in K_{\Delta_{1}},\,x_{\Delta_{2}}\in K_{\Delta_{2}},\,x_{\Delta_{12}}=u.\end{array}\right.$
Since (1.1) is feasible and $U_{0}$ is given as in (A.3), we have
(A.10) $\min_{u\in U_{0}}[F_{1}(u)+F_{2}(u)]\,\,=\,\,f_{\min}.$
For $i=1,\ldots,D$, denote the values
Let $p\coloneqq v_{1}\varphi_{1}(x_{\Delta_{12}})+\cdots+v_{D}\varphi_{D}(x_{\Delta_{12}}),$ then $p\in\mathbb{R}[x_{\Delta_{12}}]$ and
(A.11) $p(u)=\left\{\begin{array}[]{rl}-F_{1}(u)&\mathrm{~{}if~{}}u\in U_{0}\cup U_{1},\\ F_{2}(u)-f_{\min}&\mathrm{~{}if~{}}u\in U_{2}.\end{array}\right.$
Observe the relations
Since $\mathfrak{p}_{12}(K_{\Delta_{1}})=U_{0}\cup U_{1}$, (A.11) implies
[…]
Since $\mathfrak{p}_{21}(K_{\Delta_{2}})=U_{0}\cup U_{2}$, it holds
Also note that
Then (A.10) implies
$\min_{u\in U_{2}}[F_{2}(u)-p(u)-f_{\min}]\,=\,\min_{u\in U_{0}}[F_{1}(u)+F_{2}(u)-f_{\min}]=0.$
This completes the proof. ∎
In the following, we prove Theorem 4.6 by induction.
Proof of Theorem 4.6..
The conclusion holds when $m=2$, by Lemma A.1. Suppose it holds for $m=k$. We prove it also holds for $m=k+1$. The problem (1.1) can be viewed to have two sparsity blocks: $\widehat{\Delta}_{m-1}$ and $\Delta_{m}$ (see (4.19) for the notation). Consider the following two optimization problems:
(A.12) $\left\{\begin{array}[]{cl}\min&f_{1}(x_{\Delta_{1}})+\cdots+f_{m-1}(x_{\Delta_{m-1}})\\ \mbox{s.t.}&x_{\Delta_{i}}\in K_{\Delta_{i}},\,\,i=1,\ldots,m-1,\end{array}\right.\quad\left\{\begin{array}[]{cl}\min&f_{m}(x_{\Delta_{m}})\\ \mbox{s.t.}&x_{\Delta_{m}}\in K_{\Delta_{m}}.\end{array}\right.$
The feasible set for the above left optimization problem is $K_{\widehat{\Delta}_{m-1}}$ (see (4.20) for the notation). By the assumption, both $K_{\widehat{\Delta}_{m-1}}$ and $K_{\Delta_{m}}$ are nonempty and compact; the optimal values of both optimization problems in (A.12) are finite. By Lemma A.1, there exists $p\in\mathbb{R}[x_{\widehat{\Delta}_{m-1}}]\cap\mathbb{R}[x_{\Delta_{m}}]$ such that
(A.13) $\left\{\begin{array}[]{ll}f_{1}+\cdots+f_{m-1}+p\geq 0\quad\mbox{on}\quad K_{\widehat{\Delta}_{m-1}},\\ f_{m}-p-f_{\min}\geq 0\quad\mbox{on}\quad K_{\Delta_{m}}.\end{array}\right.$
$\tilde{f}_{i}\coloneqq\left\{\begin{array}[]{ll}f_{i}&\enspace\mathrm{~{}if~{}}i\neq i^{*},\\ f_{i}+p&\enspace\mathrm{~{}if~{}}i=i^{*}.\end{array}\right.$
(A.14) $\left\{\begin{gathered}\tilde{q}_{1}+\cdots+\tilde{q}_{m-1}=0,\\ \tilde{f}_{i}+\tilde{q}_{i}\geq 0~{}\text{on}~{}K_{\Delta_{i}},\,i=1,\ldots,m-1.\end{gathered}\right.$
Let $q_{m}\coloneqq-p-f_{\min}$, and for each $i=1,\ldots,m-1$, let
Then, we have $q_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ for every $i\in[m]$ and
Therefore, (A.13) and (A.14) imply that
(A.15) $\left\{\begin{gathered}q_{1}+\cdots+q_{m}+f_{\min}=0,\\ f_{i}+q_{i}\geq 0~{}\text{on}~{}K_{\Delta_{i}},\,i=1,\ldots,m.\end{gathered}\right.$
This completes the proof. ∎

### v2

@@ 6. Some examples
7. Some proofs
7.1. Proof of Theorem 3.3
Proof.
(Proof of Theorem 3.3) Since (3.8) holds, we have
$\rho\coloneqq\min_{1\leq i\leq m}\rho_{i}.$
For the case $10<\rho<$ 1, define the new tms
${V}_{h_{i}}^{spa,2t}[y^{*}]$ $=$ $\rho{V}_{h_{i}}^{spa,2t}[\hat{y}]+(1-\rho){V}_{h_{i}}^{spa,2t}[\tilde{y}],$
${V}_{h_{i}}^{spa,2t}[\tilde{y}]$ $=$ $\frac{1}{1-\rho}\left({V}_{h_{i}}^{spa,2t}[y^{*}]-\rho{V}_{h_{i}}^{spa,2t}[\hat{y}]\right)=0,$
On the other hand, it also holds that
Since $10<\rho<$ 1, we must have
$f(x^{*})=f_{t}^{smo}\leq f_{\min}\leq f(x^{*}).$
This shows that $x^{*}$ is a minimizer of (1.1). ∎
7.2. Proof of Theorem 3.5
Proof.
(Proof of Theorem 3.5) Pick arbitrary $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, then (3.15) holds, by the given assumption. The flat truncation (3.13) implies the decomposition (3.8). Note that
Up to permutation, we have
Next, we show that there exist points $u^{(1)},\ldots,u^{(r)}\in\mathbb{R}^{n}$ such that (3.16) holds, by induction on $m$. For the base step (i.e., $m=2$), let $u^{(l)}$ be such that ($l=1,\ldots,r$)
$(u^{(l)})_{k}=(u^{(1,l)})_{k}\ \text{ for }\ k\in\Delta_{1},\quad(u^{(l)})_{k}=(u^{(2,l)})_{k}\ \text{ for }\ k\in\Delta_{2}.$
Since (7.1) holds for $i=1$ and $j=2$, the above $u^{(l)}$ is well-defined. The decomposition (3.16) holds for such $u^{(l)}$. For the inductive step, assume the conclusion holds for $m-1$, and we prove it for $m$. By the induction, there exist
$\bar{u}^{(1)},\ldots,\bar{u}^{(r)}\in\mathbb{R}^{\bar{n}},\quad\text{where}\quad\bar{n}=|\Delta_{1}\cup\cdots\cup\Delta_{m-1}|,$
such that each $u^{(i,l)}$ is the projection of $\bar{u}^{(l)}$ from $\mathbb{R}^{\bar{n}}$ to $\mathbb{R}^{\Delta_{i}}$. Since $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ and satisfies the RIP, we have $\emptyset\neq\Delta_{m}\cap(\cup_{i=1}^{m-1}\Delta_{i})\subseteq\Delta_{i^{*}}$ for some $i^{*}\in[m-1]$. Then, we construct vectors $u^{(1)},\ldots,u^{(r)}\in\mathbb{R}^{n}$ such that
$(u^{(l)})_{k}=(\bar{u}^{(l)})_{k}\quad\text{for}\quad k\in\Delta_{1}\cup\cdots\cup\Delta_{m-1},$
$(u^{(l)})_{k}=(u^{(m,l)})_{k}\quad\text{for}\quad k\in\Delta_{m}.$
Since (7.1) holds for $i=i^{*}$ and $j=m$, the above $u^{(l)}$ is well-defined. Therefore, every $u^{(i,l)}$ is the projection of $u^{(l)}$ from $\mathbb{R}^{n}$ to $\mathbb{R}^{\Delta_{i}}$. This means that (3.16) holds.
Since $(u^{(j)})_{\Delta_{i}}\in K_{\Delta_{i}}$ for all $i$, we have $u^{(j)}\in K$ for all $j$. So every $[u^{(j)}]_{spa,2t}$ is feasible for (3.3) with the relaxation order equal to $t$, and every $[u^{(j)}]_{spa,2k}$ is feasible for (3.3) with the relaxation order equal to $k$. Therefore,
for all $j=1,\ldots,r$. Moreover, the decomposition (3.16) implies
Since each $\lambda_{i}>0$, the above implies that for all $j=1,\ldots,r$,
7.3. Proof of Theorem 4.4
Proof.
(Proof of Theorem 4.4) (i) Since each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$ is archimedean, the feasible sets $K_{\Delta_{i}}$ and $K$ are all compact. So, (1.1) achieves its minimum value, say, at a minimizer $u\in K$. Assumption 4.1 implies
$(f_{1}+p_{1})+\cdots+(f_{m}+p_{m})=f-f_{\min}.$
Since $f(u)=f_{\min}$ and each $f_{i}+p_{i}\geq 0$ on $K_{\Delta_{i}}$, the minimum value of (4.6) is 00. By [31, Theorem 1.1], if the LICQC, SCC, and SOSC hold at every minimizer of (4.6), then there is a relaxation order $N_{i}$ such that for all $\epsilon>0$,
Let $N\coloneqq\max\{N_{1},\ldots,N_{m}\}$. The above then implies
By Assumption 4.1, we have $p_{1}+\dots+p_{m}+f_{\min}=0$, so for all $k\geq N$
$\in$ $\operatorname{Ideal}[h]_{spa,k}+\operatorname{QM}[g]_{spa,k}.$
This follows from [25, Theorem 9.5.3] and [31, Theorem 3.1]. So, we get
By Assumption 4.1, $p_{1}+\cdots+p_{m}+f_{\min}=0$, so
(7.4) $\left\{\begin{array}[]{cll}\max&\gamma\\ \mathrm{s.t.}&f_{i}+p_{i}-\gamma\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k}.\end{array}\right.$
Its dual optimization is the moment relaxation
(7.5) $\left\{\begin{array}[]{cll}\min&\langle f_{i}+p_{i},y_{\Delta_{i}}\rangle\\ \mathrm{s.t.}&{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]=0,L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}]\succeq 0,\\ &(y_{\Delta_{i}})_{0}=1,\ M_{\Delta_{i}}^{(k)}[y_{\Delta_{i}}]\succeq 0,\\ &y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}^{\Delta_{i}}_{2k}}.\end{array}\right.$
7.4. Proof of Theorem 4.5
Proof.
(Proof of Theorem 4.5) (i) Since each $V_{\mathbb{R}}(h_{i})$ is finite, there exists $N_{i}$ such that for all $\epsilon>0$,
$\in$ $\operatorname{Ideal}[h]_{spa,2N}+\operatorname{QM}[g]_{spa,2N}.$
$f_{\min}-m\epsilon\leq f_{N}^{spa}\leq f_{\min}.$
This forces $f_{k}^{spa}=f_{\min}$ for all $k\geq N$.
7.5. Proof of Theorem 4.6
We first prove Theorem 4.6 for the case $m=2$.
Lemma 7.1.
(7.6) $\left\{\begin{array}[]{cll}\displaystyle\min&f_{1}(x_{\Delta_{1}})\\ \mathrm{s.t.}&h_{1}(x_{\Delta_{1}})=0,\\ &g_{1}(x_{\Delta_{1}})\geq 0,\end{array}\right.\quad\left\{\begin{array}[]{cll}\displaystyle\min&f_{2}(x_{\Delta_{2}})\\ \mathrm{s.t.}&h_{2}(x_{\Delta_{2}})=0,\\ &g_{2}(x_{\Delta_{2}})\geq 0.\end{array}\right.$
(7.7) $f_{1}+p\geq 0\,\ \mathrm{on}\,\ K_{\Delta_{1}},\quad f_{2}-p-f_{\min}\geq 0\,\ \mathrm{on}\,\ K_{\Delta_{2}},$
Proof.
By the assumption, we can write that
(7.8) $U_{0}\coloneqq\mathfrak{p}_{12}(K_{\Delta_{1}})\cap\mathfrak{p}_{21}(K_{\Delta_{2}}),\enspace U_{1}\coloneqq\mathfrak{p}_{12}(K_{\Delta_{1}})\backslash U_{0},\enspace U_{2}\coloneqq\mathfrak{p}_{21}(K_{\Delta_{2}})\backslash U_{0}.$
Define the optimal value functions:
(7.11) $\left\{\begin{array}[]{rcl}F_{1}(u)\coloneqq&\displaystyle\min&f_{1}(x_{\Delta_{1}})\\ &\mathrm{s.t.}&x_{\Delta_{1}}\in K_{\Delta_{1}},\,x_{\Delta_{12}}=u,\end{array}\right.$
(7.14) $\left\{\begin{array}[]{rcl}F_{2}(u)\coloneqq&\displaystyle\min&f_{2}(x_{\Delta_{2}})\\ &\mathrm{s.t.}&x_{\Delta_{2}}\in K_{\Delta_{2}},\,x_{\Delta_{12}}=u.\end{array}\right.$
For each $u\in U_{0}\cup U_{1}$, the problem (7.11) is feasible and its optimal value is finite. Similarly, for each $u\in U_{0}\cup U_{2}$, the problem (7.14) is feasible and its optimal value is finite. For all $u\in U_{0}$, $F_{1}(u)+F_{2}(u)$ equals the minimum value of
$\left\{\begin{array}[]{cll}\displaystyle\min&f_{1}(x_{\Delta_{1}})+f_{2}(x_{\Delta_{2}})\\ \mathrm{s.t.}&x_{\Delta_{1}}\in K_{\Delta_{1}},\,x_{\Delta_{2}}\in K_{\Delta_{2}},\,x_{\Delta_{12}}=u.\end{array}\right.$
Since (1.1) is feasible and $U_{0}$ is given as in (7.8), we have
(7.15) $\min_{u\in U_{0}}[F_{1}(u)+F_{2}(u)]\,\,=\,\,f_{\min}.$
For $i=1,\ldots,D$, denote the values
Let $p\coloneqq v_{1}\varphi_{1}+\cdots+v_{D}\varphi_{D}$, then $p\in\mathbb{R}[x_{\Delta_{12}}]$ and
(7.16) $p(u)=\left\{\begin{array}[]{rl}-F_{1}(u)&\mathrm{~{}if~{}}u\in U_{0}\cup U_{1},\\ F_{2}(u)-f_{\min}&\mathrm{~{}if~{}}u\in U_{2}.\end{array}\right.$
Observe the relations
Since $\mathfrak{p}_{12}(K_{\Delta_{1}})=U_{0}\cup U_{1}$, (7.16) implies
[…]
Since $\mathfrak{p}_{21}(K_{\Delta_{2}})=U_{0}\cup U_{2}$, it holds
Also note that
So, $\min\limits_{u\in U_{2}}[F_{2}(u)-p(u)-f_{\min}]=0$, and (7.15) implies
$\min_{u\in U_{0}}[F_{2}(u)-p(u)-f_{\min}]\,=\,\min_{u\in U_{0}}[F_{1}(u)+F_{2}(u)-f_{\min}]=0.$
This completes the proof. ∎
In the following, we prove Theorem 4.6 by induction.
Proof of Theorem 4.6..
The conclusion holds when $m=2$, by Lemma 7.1. Suppose it holds for $m=k$. We prove it also holds for $m=k+1$. The problem (1.1) can be viewed to have two sparsity blocks: $\widehat{\Delta}_{m-1}$ and $\Delta_{m}$ (see (4.15) for the notation). Consider the following two optimization problems:
(7.17) $\left\{\begin{array}[]{cl}\min&f_{1}(x_{\Delta_{1}})+\cdots+f_{m-1}(x_{\Delta_{m-1}})\\ \mathrm{s.t.}&x_{\Delta_{i}}\in K_{\Delta_{i}},\,\,i=1,\ldots,m-1,\end{array}\right.\quad\left\{\begin{array}[]{cl}\min&f_{m}(x_{\Delta_{m}})\\ \mathrm{s.t.}&x_{\Delta_{m}}\in K_{\Delta_{m}}.\end{array}\right.$
The feasible set for the above left optimization problem is $K_{\widehat{\Delta}_{m-1}}$ (see (4.16) for the notation). By the assumption, both $K_{\widehat{\Delta}_{m-1}}$ and $K_{\Delta_{m}}$ are nonempty and compact; the optimal values of both optimization problems in (7.17) are finite. By Lemma 7.1, there exists $p\in\mathbb{R}[x_{\widehat{\Delta}_{m-1}}]\cap\mathbb{R}[x_{\Delta_{m}}]$ such that
(7.18) $\left\{\begin{array}[]{ll}f_{1}+\cdots+f_{m-1}+p\geq 0\quad\mbox{on}\quad K_{\widehat{\Delta}_{m-1}},\\ f_{m}-p-f_{\min}\geq 0\quad\mbox{on}\quad K_{\Delta_{m}}.\end{array}\right.$
$\tilde{f}_{i}\coloneqq\left\{\begin{array}[]{ll}f_{i},&\enspace\mathrm{~{}if~{}}i\neq i^{*},\\ f_{i}+p,&\enspace\mathrm{~{}if~{}}i=i^{*}.\end{array}\right.$
(7.19) $\left\{\begin{gathered}\tilde{q}_{1}+\cdots+\tilde{q}_{m-1}=0,\\ \tilde{f}_{i}+\tilde{q}_{i}\geq 0~{}\text{on}~{}K_{\Delta_{i}},\,i=1,\ldots,m-1.\end{gathered}\right.$
Let $q_{m}\coloneqq-p-f_{\min}$, and for each $i=1,\ldots,m-1$, let
Then, we have $q_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ for every $i\in[m]$ and
Therefore, (7.18) and (7.19) imply that
(7.20) $\left\{\begin{gathered}q_{1}+\cdots+q_{m}+f_{\min}=0,\\ f_{i}+q_{i}\geq 0~{}\text{on}~{}K_{\Delta_{i}},\,i=1,\ldots,m.\end{gathered}\right.$
This completes the proof. ∎
7.6. Proof of Theorem 5.2
Proof.
(Proof of Theorem 5.2) (i) As we have shown in Theorem 4.4(i), the minimum value of (4.6) is 00. Since $K_{\Delta_{i}}$ is finite, there exists a degree $N_{i}$ such that for all $\epsilon>0$,
This is shown in the proof of Theorem 4.1 of [30]. Note that
$\in$ $\operatorname{Ideal}[h]_{spa,2N}+\operatorname{Pre}[g]_{spa,2N}.$
$f_{\min}-m\epsilon\leq f_{N}^{smg}\leq f_{\min}.$
This forces $f_{N}^{smg}=f_{\min}$, so $f_{k}^{smg}=f_{\min}$ for all $k\geq N$.
(7.21) $\left\{\begin{array}[]{cll}\min&\langle f_{i}+p_{i},y_{\Delta_{i}}\rangle\\ \mathrm{s.t.}&{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]=0\,(i\in[m]),\\ &L_{g_{i,J}}^{\Delta_{i},k}[y_{\Delta_{i}}]\succeq 0\,(i\in[m],\,J\subseteq[s_{i}]),\\ &(y_{\Delta_{i}})_{0}=1,\ y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}^{\Delta_{i}}_{2k}}.\end{array}\right.$
Since each $K_{\Delta_{i}}$ is finite, by [32, Theorem 5.6.7], we know $y^{*}_{\Delta_{i}}$ satisfies the flat truncation (5.5) when $k$ is large enough. The remaining part follows from Theorem 3.3.
(iii) Note that $\mathcal{X}_{\Delta_{i}}\subseteq K_{\Delta_{i}}$, so $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$ is also a finite set. Then, by the same proof as for Theorem 4.4(iii), we can get (5.6) and (5.7). ∎
@@ Acknowledgements
This project was begun at a SQuaRE at the American Institute of Mathematics (AIM). The authors would like to thank AIM for providing a supportive and mathematically rich environment. Jiawang Nie and Linghao Zhang are partially supported by the National Science Foundation grant DMS-2110780. Zheng Qu is partially supported by the Research Center for Intelligent Operations Research at Department of Applied Mathematics, the Hong Kong Polytechnic University. Xindong Tang is partially supported by the Hong Kong Research Grants Council HKBU-15303423.